Deepesh:

A manufacturer is designing a two-wheeled cart that can maneuver through tight spaces. On one test model, the wheel placement (center) and radius is modeled by the equation (x + 2) 2 + (y − 1) 2 = 4. What is the graph that shows the position and radius of the wheels?

Nnesha:

the equation of a circle \[\large\rm (x-h)^2+(y-k)^2=r^2\] where `(h,k)` is the center point and `r` represent the radius.

Nnesha:

\[\large\rm (x-\color{red}{h})^2+(y-\color{blue}{k})^2=\color{orange}{r}^2\]\[\large\rm (x+\color{red}{2})^2+(y-\color{blue}{1})^2=\color{orange}{4}\] we can rewrite 4 as 2^2 and (x+2)^2 is same as (x-(-2))^2